# Tensor-Var: Variational Data Assimilation in Tensor Product Feature Space

### v2

@@ 2 Background and Problem Formulation
2.1 4D variational data assimilation
Consider a dynamical system in discrete-time comprising dynamical model and observation model:
in which $F$ is the dynamical model that advances the state $s_{t}$ to $s_{t+1}$, and $G$ is the observation model that maps the $s_{t}$ to the observation $o_{t}$. The noise components $\epsilon^{s}_{t},\epsilon^{o}_{t}$ are assumed to follow the zero mean Gaussian distributions with covariance matrices $Q$ and $R$. The objective of (weakly constraint) 4D-variational DA is to minimize a cost function:
$J(s_{0:T})=\|s_{0}-s^{b}_{0}\|^{2}_{B^{-1}}+\sum_{t=0}^{T}\|o_{t}-G(s_{t})\|^{2}_{R^{-1}}+\sum_{t=1}^{T}\|s_{t}-F(s_{t-1})\|^{2}_{Q^{-1}},$ (2)
in which $s_{0}^{b}$ is a prior guess for the initial state $s_{0}$, with $B$ as the background covariance matrix representing the uncertainty, i.e., $s_{0}\sim\mathcal{N}(s_{0}^{b},B)$. The second and third terms account for errors in the observation and dynamical models in equation 1.
In systems with nonlinear dynamics and observations, the cost function is typically non-convex with costly model evaluations, raising challenges in minimization of equation 2. Thus, we seek a space with a linear structure, in which the 4D-Var optimization can be efficiently solved. Kernel methods provide a framework for linearization by projecting data into an infinite-dimensional feature space (Jacot et al., 2018; Bevanda et al., 2024).
2.2 Kernel conditional mean embedding
A Reproducing Kernel Hilbert Space (RKHS) $\mathbb{H}_{S}$ with kernel $k_{S}$ is a Hilbert space, satisfying the reproducing property (Schölkopf & Smola, 2002). Let $k_{S}$ and $k_{O}$ be positive-definite (pd) kernels on RKHSs $\mathbb{H}_{S}$ and $\mathbb{H}_{O}$. We denote the kernel features as $\phi_{O}(o)=k_{O}(o,\cdot)$ and $\phi_{S}(s)=k_{S}(s,\cdot)$, referring the $\mathbb{H}_{S}$ and $\mathbb{H}_{O}$ as feature spaces. In this paper, we use the word kernel feature and the word feature interchangeably.
The kernel mean embedding of a distribution of $S$ is defined as the expectation of feature, $\mathbb{E}[\phi_{S}(S)]\in\mathbb{H}_{S}$ (Fukumizu et al., 2011). For characteristic kernels, these embeddings are injective, uniquely determining the probability distribution (Muandet et al., 2017). In addition to mean embedding, we will need the (uncentered) covariance operator (Baker, 1973) defined as $\mathcal{C}_{SO}=\mathbb{E}[\phi_{S}(S)\otimes\phi_{O}(O)]$, where $\otimes$ denotes the tensor product. This operator is also the embedding of the joint distribution as an element in the tensor product Hilbert space $\mathbb{H}_{S}\otimes\mathbb{H}_{O}$. The covariance operators extend the concepts of covariance matrices from finite dimensional spaces to infinite feature spaces.
Figure 1: Demonstration of Tensor-Var: A DA system with nonlinear dynamical model, observation model, and non-convex cost function (bottom) can be represented linearly in feature space using kernel conditional mean embeddings that results in a convex cost function (Jacot et al., 2018) (top).
Conditional mean embedding.
To represent the dynamical and observation models in equation 1, the conditional mean embedding (CME) plays an important role. The CME of a conditional distribution of $S$ given $O=o$ is defined as the conditional expectation of kernel features $\mathbb{E}[\phi_{S}(S)|O=o]$. Under standard assumptions111(1) $\mathcal{C}_{SS}$ is injective, and (2) $\mathbb{E}[f(S)|O=o]\in\mathbb{H}_{S}$ for any $f\in\mathbb{H}_{S}$ and $o\in\mathbb{R}^{n_{o}}$., there exists a linear operator $\mathcal{C}_{S|O}\colon\mathbb{H}_{O}\rightarrow\mathbb{H}_{S}$ such that $\mathbb{E}[\phi_{S}(S)|O=o]=\mathcal{C}_{S|O}\phi_{O}(o)$. Given independent and identically distributed samples $\{(s_{i},o_{i})\}_{i=1}^{N}$, an empirical estimate of the operator can be obtained as follows:
$\hat{\mathcal{C}}_{S|O}=\hat{\mathcal{C}}_{SO}(\hat{\mathcal{C}}_{OO}+\lambda I)^{-1},$ (3)
where $\lambda$ is the regularization parameter, the $\hat{\mathcal{C}}_{SO}$ and $\hat{\mathcal{C}}_{OO}$ are empirical covariance operators, e.g. $\hat{\mathcal{C}}_{SO}=\frac{1}{N}\sum_{i=1}^{N}\phi_{S}(s_{i})\otimes\phi_{O}(o_{i})$, and $I$ denotes the identity matrix. Alternatively, the empirical estimate $\hat{\mathcal{C}}_{S|O}$ is equivalently to the minimizer of the following regression problem,
$\hat{L}(\mathcal{C})=\frac{1}{N}\sum_{i=1}^{N}\|\phi_{S}(s_{i})-\mathcal{C}\phi_{O}(o_{i})\|^{2}+\lambda\|\mathcal{C}\|^{2},$ (4)
which offers a way to linearize the DA systems over the feature spaces.
3 Methods
In this section, we introduce our Tensor-Var approach, which embeds 4D-Var into the kernel feature space and provides a theoretical analysis demonstrating the existence of linear dynamics with consistent convergence between the original and feature space solutions. To effectively address incomplete observations, we propose an inverse observation operator that leverages consecutive historical observations. Finally, we also propose a method to learn adaptive deep features (DFs) using neural networks within the Tensor-Var framework, improving real-world applicability.
3.1 CME of 4D-Var in RKHS
Having introduced the necessary tools for manipulating kernel embeddings, we now focus on learning the linearized models of the system in equation 1.
CME of dynamical model.
Let $S^{+}$ be the one-step forward of $S$. Given the dynamical model $F$ in equation 1 and kernel feature $\phi_{S}$, the CME operator $\mathcal{C}_{S^{+}|S}$ can be recognized as the best linear approximation in the feature space $\mathbb{H}_{S}$ that minimize the regression residual $\mathbb{E}\big{[}\|\phi_{S}(S^{+})-\mathcal{C}\phi_{S}(S)\|^{2}\big{]}$. Given finite data $\{(s_{i}^{+},s_{i})\}_{i=1}^{N}$ sliced from the system trajectory $s_{1:N+1}$, we can obtain the empirical estimate $\hat{\mathcal{C}}_{S^{+}|S}$ as equation 3 with theoretical supports of convergence (Fukumizu et al., 2013; Klus et al., 2020). The CME operator $\hat{\mathcal{C}}_{S^{+}|S}$ effectively characterizes the system dynamics as a linear model and simplifies the 4D-Var as a convex optimization in the feature space $\mathbb{H}_{S}$.
CME of inverse observation model.
Analogous to the dynamical model, the observation model $G$ in equation 1 can be linearized by the CME operator $\mathcal{C}_{O|S}$, which has been used as observation models for filtering algorithms (Song et al., 2009; Fukumizu et al., 2013; Kanagawa et al., 2016; Gebhardt et al., 2019). Most approaches assume a complete observation setting, where the observations can fully determine the state. In practice, observations are often incomplete representation of state, with $n_{s}>n_{o}$, leading to underdetermined systems (Liu et al., 2022). In such systems, the lack of a bijective mapping between the state and observation spaces means that observations cannot uniquely determine the system’s state. As a result, the optimization problem in 4D-Var may produce sub-optimal solutions (Asch et al., 2016). It is information-theoretically impossible to distinguish any two mixtures of states based on a single-step observation if $n_{s}>n_{o}$. By introducing the past $m$ consecutive observations as history $h_{t}=o_{t-m-1:t-1}\in\mathbb{R}^{m\times n_{o}}$, the joint information from history and current observation is enough to estimate the system state. The choice of history length is critical: too short lacks sufficient information, while too long is inefficient. Empirically, we performed an ablation study to assess the effects of history length, as detailed in the subsection 4.4.
To effectively incorporate history, we introduce another kernel feature $\phi_{H}$ with an induced RKHS $\mathbb{H}_{H}$. The space $\mathbb{H}_{OH}=\mathbb{H}_{O}\otimes\mathbb{H}_{H}$ is called a tensor product RKHS on $O\times H$ with associated kernel feature $\phi_{OH}(o,h)=\phi_{O}(o)\otimes\phi_{H}(h)$. As shown in (Song et al., 2013; Muandet et al., 2017), the CME operator can be extended to high-order features, allowing us to embed the joint distributions over $s_{t}$, $o_{t}$, and $h_{t}$ into the feature space (Song et al., 2009; 2013). The tensor product feature $\phi_{OH}$ captures the high-order dependencies between observations and history. The CME operator $\mathcal{C}_{S|OH}$ is the linear inverse observation model that minimizes the state estimation error given observations and history. Given dataset $\{(s_{i},o_{i},h_{i})\}_{i=1}^{N}$, the empirical counterpart $\hat{\mathcal{C}}_{S|OH}=\hat{\mathcal{C}}_{SOH}(\hat{\mathcal{C}}_{(OH)(OH)}+\lambda I)^{-1}$ follows the same way as equation 3. The $\hat{\mathcal{C}}_{SOH}$ is the empirical high order tensor, e.g., $\hat{\mathcal{C}}_{SOH}=\sum_{i=1}^{N}\phi_{S}(s_{i})\otimes\phi_{O}(o_{i})\otimes\phi_{H}(h_{i})$.
Our approach is closely related to delay embeddings, a well-established method in dynamical system theory for reconstructing attractors and state spaces from sequential data (Sauer et al., 1991; Krämer et al., 2021). As shown by Takens’ embedding theorem (Takens, 2006), the dynamics of a system can be reconstructed in a higher-dimensional space using historical observations. Parallel to Takens’ embedding theorem, recent theoretical advancements (Uehara et al., 2022) have focused on partial observability from a learning theory perspective. Liu et al. (2022) showed that the required history length for effective state reconstruction is determined by the complexity of the dynamical system.
Feature space 4D-Var.
Using the kernel features, we linearize the original nonlinear dynamics and observations in the feature space $\mathbb{H}_{S}$. This transformation enables us to reformulate the 4D-Var optimization objective equation 2 into the feature space and optimize over a sequence of elements $z_{0:T}$ in feature space $\mathbb{H}_{S}$:
$\begin{split}&\min_{z_{0:T}}\,\,\|z_{0}-\phi_{S}(s_{0}^{b})\|^{2}_{\mathcal{B}^{-1}}+\sum_{t=0}^{T}\|z_{t}-\mathcal{C}_{S|OH}\phi_{OH}(o_{t},h_{t})\|^{2}_{\mathcal{R}^{-1}}+\sum_{t=0}^{T-1}\|z_{t+1}-\mathcal{C}_{S^{+}|S}z_{t}\|^{2}_{\mathcal{Q}^{-1}},\end{split}$ (5)
where the $\mathcal{B}$, $\mathcal{R}$, and $\mathcal{Q}$ are the covariance operators for the background error, observation error, and model error in the feature space. In this work, we estimated the three operators as the empirical error covariance matrices from the training dataset (the explicit estimation can be found in Appendix B.1). As a result, our approach linearizes the original nonlinear dynamical and observation models, reducing the problem to solving a linear 4D-Var problem, with a quadratic cost function with linear dynamics $z_{t+1}=\mathcal{C}_{S^{+}|S}z_{t}$. We present the pseudo-algorithms for CME of 4D-Var and performing Tensor-Var, detailed in Algorithms 2 and 3 in Appendix B.2.
3.2 Learning the deep features within Tensor-Var
Using the pre-determined kernel features has theoretical guarantees. However, it maps data into an infinite-dimensional feature space, e.g., radial basis function kernel, making empirical estimation challenging due to polynomial scaling with sample size. On the other hand, these feature maps struggle with irregular or high-dimensional data, often resulting in poor performances. Learned deep features (DFs) have emerged as alternatives to generic kernel features (Xu et al., 2022; Kostic et al., 2023; Shimizu et al., 2024), project the data into a fixed-dimensional feature space, similar to methods like low-rank approximation (Williams & Seeger, 2000). To improve the scalability, we integrate DFs with the Tensor-Var framework and validate their effectiveness through experiments.
Learning the state feature.
Recall that the CME operator $\mathcal{C}_{S^{+}|S}$ is the best linear approximation of the system dynamics in the feature space. We propose to jointly learn the feature $\phi_{\theta_{S}}\colon\mathbb{R}^{n_{s}}\rightarrow\mathbb{R}^{d_{s}}$ with $\hat{\mathcal{\mathcal{C}}}_{S^{+}|S}$ by minimizing the loss $L(\theta_{S})=\min_{\mathcal{C}\in\mathbb{R}^{d_{s}\times d_{s}}}L(\mathcal{C},\theta_{S})=\mathbb{E}\big{[}\|\phi_{\theta_{S}}(s^{+})-\hat{\mathcal{C}}_{S^{+}|S}\phi_{\theta_{S}}(s)\|^{2}\big{]}$. Predictions using $\hat{\mathcal{C}}_{S^{+}|S}$ are in the feature space; however, for DA problems, reconstruction to the original state space is required, which is known as the preimage problem (Honeine & Richard, 2011). Here, we learn an inverse feature $\phi^{\dagger}_{\theta^{\prime}_{S}}$ to solve the preimage problem during training, avoiding repeated optimization whenever computing preimages. The final training loss is the combination of the two terms as
$\begin{split}L(\theta_{S},\theta_{S}^{\prime})&=\mathbb{E}\big{[}\|\phi_{\theta_{S}}(s^{+})-\hat{\mathcal{C}}_{S^{+}|S}\phi_{\theta_{S}}(s)\|^{2}\big{]}+w\,\mathbb{E}\Big{[}\|s-\phi^{\dagger}_{\theta^{\prime}_{S}}\big{(}\phi_{\theta_{S}}(s)\big{)}\|^{2}\Big{]},\end{split}$
where $w\in[0,1]$ is a weighting coefficient, and $\hat{\mathcal{C}}_{S^{+}|S}$ is computed as the CME over training batches. Note that using the DFs corresponds to a linear kernel in the learned feature space, where $k(s_{i},s_{j})=\phi_{\theta_{S}}(s_{i})^{T}\phi_{\theta_{S}}(s_{j})$.
Learning the observation and history features.
Similar to learning the state feature, $\mathcal{C}_{S|OH}$ is the minimizer of regression problem mapping the tensor product of observation and history features to the state feature. In this phase, we learn the DFs for observation $\phi_{\theta_{O}}\colon\mathbb{R}^{n_{o}}\rightarrow\mathbb{R}^{d_{o}}$, history $\phi_{\theta_{H}}\colon\mathbb{R}^{n_{h}}\rightarrow\mathbb{R}^{d_{h}}$, and $\mathcal{C}_{S|OH}$ jointly with the loss function:
where $\hat{\mathcal{C}}_{S|OH}$ is computed similarly over training batches in parallel, with the $\otimes$ denotes Kronecker product in practice. The learning procedure of DFs is summarized in Algorithm 1.
3.3 Theoretical analysis.
In Section 3.1, we discuss how the dynamical system $F$ in equation 1 can be embedded as a linear system in the feature space. However, two important questions remain: 1) Does such linear dynamical system exist? 2) Are the solutions of original and feature space 4D-Var consistent? In this section, we provide affirmative answers to both questions using the theory of Kazantzis-Kravaris/Luenberger (KKL) observers (Andrieu & Praly, 2006). We give a road-map of the theoretical analysis with main result and refer readers to Appendix A for the details.
Under the mild assumptions that (1) the dynamical model $F$ is first-order differentiable and (2) the kernel features are all first-order differentiable, the kernel feature $\phi_{S}$ satisfies the necessary conditions as the state transformation in the KKL observer framework. This transformation enables us to represent the nonlinear dynamical system as a linear system in a higher-dimensional feature space, as established by KKL observer theory (Tran & Bernard, 2023). This result confirms the existence of such a linear system, thus answering question 1). The KKL observer theory provides a theoretical foundation for our approach, bridging nonlinear dynamics and linear 4D-Var methods (Andrieu & Praly, 2006). A detailed derivation proving that $\phi_{S}$ satisfies the conditions as the state transformation of the KKL observer can be found in Appendix A.
We consider the nonlinear system in equation 1 within a compact state space and assume that the cost function in 2 has a unique solution. Given that $\phi_{S}$ is a state transformation in the KKL observer, the system in the original state space can be represented linearly in the feature space. The solution in the feature space have consistent convergence to the unique solution of original 4D-Var problem, minimizing with respect to the cost function 5, answering the question 2). A formal theorem with detailed proofs can be found in Theorem A.4 in Appendix A.2.
4 Numerical Experiments
@@
To evaluate our proposed method, the comparison is conducted on a series of benchmark domains, representing optimization problem equation 2 of increasing complexity, including (1) the Lorenz 96 system (Lorenz, 1996) with 40 and 80 state dimensions. (2) The Kuramoto-Sivashinsky (KS) equation: a fourth order nonlinear PDE system (Papageorgiou & Smyrlis, 1991) with 128 and 256 dimensions, representing different spatial resolution. For both systems, we use a nonlinear observation model $o=G(s)=5\text{arctan}(s\pi/10)+\epsilon$, where $\epsilon$ is white noise with a standard deviation of $0.01$ times the standard deviation of state variable distribution and only $20\%$ states can be observed. To assess the practical applicability of Tensor-Var, we evaluate its performance in global medium-range weather forecasting (i.e., 3-5 days) by using a subset of the ECMWF Reanalysis v5 (ERA5) dataset for training and testing, with further details in subsection 4.2 and (Rasp et al., 2024). Furthermore, we incorporate observations extracted from real-time weather satellites track into the NWP experiment with higher spatial-resolution, showcasing the practical utility of Tensor-Var in subsection 4.3.
@@
We compare our method against several baseline approaches: (1) 3D-Var with a known observation model, (2) a model-based 4D-Var algorithm that assumes known dynamical and observation models, (3) a learned inverse observation model with the known dynamical model, as proposed by Frerix et al. (2021), and (4) In the two NWP problems, we include two more baselines: Latent 3D-Var and Latent 4D-Var (Cheng et al., 2024), which perform variational DA in a compressed latent space learned via autoencoder, with Latent 4D-Var additionally learning the dynamics in the latent space. These competitive baselines cover both operational Var-DA and ML-hybrid Var-DA methods.
Table 1: Comparison of DA performances. All baseline methods use the strong-constraint 4D-Var objective, while our approach uses the weak-constraint 4D-Var objective. Evaluation times are reported as each assimilation window’s mean and standard deviation. Our method consistently outperforms the baselines across all benchmark domains.
[…]
3D-Var $14.17\pm 0.93$ $12.59\mathbf{\pm 0.39}$
Lorenz 96 4D-Var $12.27\pm 1.41$ $210.52\pm 3.87$
$n_{s}=40,n_{o}=8$ Frerix et al. (2021) $9.89\pm 1.63$ $167.43\pm 1.33$
Ours $\mathbf{8.32\pm 0.87}$ $\mathbf{12.51}\pm 1.97$
3D-Var $15.19\pm 1.09$ $\mathbf{19.38\pm 0.37}$
Lorenz 96 4D-Var $12.38\pm 1.11$ $322.21\pm 5.73$
$n_{s}=80,n_{o}=16$ Frerix et al. (2021) $10.79\pm\mathbf{0.57}$ $286.11\pm 2.43$
Ours $\mathbf{9.04}\pm 1.32$ $21.09\pm 0.79$
3D-Var $17.64\pm 1.27$ $\mathbf{16.48\pm 1.17}$
Kuramoto-Sivashinsky 4D-Var $15.46\pm 1.07$ $94.83\pm 3.89$
$n_{s}=128,n_{o}=32$ Frerix et al. (2021) $10.25\pm 1.34$ $63.28\pm 1.91$
Ours $\mathbf{9.69}\pm 1.56$ $19.58\pm 1.23$
3D-Var $16.66\pm 0.69$ $\mathbf{15.81\pm 0.92}$
Kuramoto-Sivashinsky 4D-Var $10.67\pm 0.62$ $95.68\pm 1.35$
$n_{s}=256,n_{o}=64$ Frerix et al. (2021) $8.87\pm 0.55$ $68.39\pm 1.23$
Ours $\mathbf{4.31\pm 0.19}$ $17.37\pm 1.36$
@@
In each experiment, we measure the quality of assimilation using the Normalized Root Mean Square Error (NRMSE) $\sqrt{\frac{1}{T}\sum_{t=0}^{T}\|\hat{s}_{t}-s_{t}\|^{2}}/(s_{\text{max}}-s_{\text{min}})$ over the assimilation window, where $s_{\text{max}},s_{\text{min}}$ are the maximum and minimum state value in training dataset. These results and the average evaluation time for each algorithm are reported in Table 1. All metrics are evaluated 20 times with different initial conditions, reporting the mean and standard deviation. For Tensor-Var, the history length $m$ is selected using a cross-validation approach as an ablation study in subsection 4.4, and the objective function in feature space is minimized using quadratic programming methods, implemented via CVXPY (Diamond & Boyd, 2016). Baselines use the L-BFGS method (Nocedal & Wright, 2006) with 10 history vectors for the Hessian approximation. For 4D-Var baselines, the adjoint method is not used, and the strongly constrained 4D-Var is applied. The background state $s_{b}$ is set to the average state of the training set.
As shown in Table 1, our method consistently outperforms the other baseline methods in all metrics. In all four tasks, Tensor-Var achieves the lowest mean and standard deviation of NRMSE for assimilation accuracy, demonstrating strong generalization from the Lorenz-96 systems to the more complex KS systems. The better performances are attributed to the linearization of the dynamics and the history-augmented inverse observation operator, which makes the 4D-Var optimization convex with more reliable convergence to the global optimum in the feature space. The larger NRMSE errors observed in the model-based 3D-Var and 4D-Var methods are due to that incomplete observations lead to the uncontrolled errors in the unobserved state dimensions. Although the ML-hybrid 4D-Var method (Frerix et al., 2021) employs a learned inverse observation model, it struggles with generalization due to the incomplete observations, leading to poor performance on test data. By incorporating historical information, our approach more effectively controls estimation errors in the unobserved state dimensions. This improvement is clearly demonstrated in Figures 8 and 9 in Appendix C.1.2 and C.1.2, which qualitatively compares assimilation performance in the Lorenz-96 and KS systems. In terms of computational efficiency, the linearity of Tensor-Var makes it faster than all 4D-Var baselines while also outperforming or being comparable to the static 3D-Var method, attributing to the reduced complexity of solving convex optimization problems in the feature space.

### v3

@@ 4 Experiment
3D-Var $14.17\pm 0.93$ $12.59\mathbf{\pm 0.39}$ $13\pm 2$
Lorenz 96 4D-Var $12.27\pm 1.41$ $210.52\pm 3.87$ $17\pm 6$
$n_{s}=40$ 4D-Var by Adjoint $12.18\pm 1.46$ $33.78\pm 2.11$ $16\pm 4$
$n_{o}=8$ Frerix et al. (2021) $9.89\pm 1.63$ $167.43\pm 1.33$ $11\pm 3$
Ours $\mathbf{8.32\pm 0.87}$ $\mathbf{12.51}\pm 1.97$ $\mathbf{8}\pm\mathbf{2}$
3D-Var $15.19\pm 1.09$ $\mathbf{19.38\pm 0.37}$ $14\pm 2$
Lorenz 96 4D-Var $12.38\pm 1.11$ $322.21\pm 5.73$ $22\pm 8$
$n_{s}=80$ 4D-Var by Adjoint $12.44\pm 2.55$ $48.07\pm 2.59$ $22\pm 8$
$n_{o}=16$ Frerix et al. (2021) $10.79\pm\mathbf{0.57}$ $286.11\pm 2.43$ $14\pm 2$
Ours $\mathbf{9.04}\pm 1.32$ $21.09\pm 0.79$ $\mathbf{9}\pm\mathbf{1}$
3D-Var $17.64\pm 1.27$ $\mathbf{16.48\pm 1.17}$ $\mathbf{13}\pm\mathbf{0}$
Kuramoto-Sivashinsky 4D-Var $15.46\pm 1.07$ $94.83\pm 3.89$ $77\pm 4$
$n_{s}=128$ 4D-Var by Adjoint $15.43\pm\mathbf{0.87}$ $22.14\pm 4.01$ $72\pm 4$
$n_{o}=32$ Frerix et al. (2021) $10.25\pm 1.34$ $63.28\pm 1.91$ $28\pm 1$
Ours $\mathbf{9.69}\pm 1.56$ $19.58\pm 1.23$ $18\pm 2$
3D-Var $16.66\pm 0.69$ $\mathbf{18.81\pm 0.92}$ $\mathbf{15}\pm\mathbf{1}$
Kuramoto-Sivashinsky 4D-Var $10.67\pm 0.62$ $99.68\pm 2.35$ $83\pm 3$
$n_{s}=256$ 4D-Var by Adjoint $10.23\pm 0.21$ $25.77\pm 2.13$ $81\pm 3$
$n_{o}=64$ Frerix et al. (2021) $8.87\pm 0.55$ $71.39\pm 1.23$ $31\pm 2$
Ours $\mathbf{4.31\pm 0.19}$ $21.37\pm 1.36$ $23\pm 3$
@@
To evaluate our proposed method, the comparison is conducted on a series of benchmark domains, representing the optimization problem (2) of increasing complexity, including (1) Lorenz 96 system (Lorenz, 1996) with $n_{s}=40$ and $80$. (2) Kuramoto-Sivashinsky (KS) equation: a fourth-order nonlinear PDE system (Papageorgiou & Smyrlis, 1991) with $n_{s}=128$ and $256$, representing different spatial resolution. For both systems, we use a nonlinear observation model $o=G(s)=5\text{arctan}(s\pi/10)+\epsilon$, where $\epsilon$ is white noise with a standard deviation of $0.01$ times the standard deviation of the state variable distribution, and only $20\%$ states can be observed. To assess the practical applicability of Tensor-Var, we evaluate its performance in global medium-range weather forecasting (i.e. 3-5 days) by using a subset of the ECMWF Reanalysis v5 (ERA5) dataset for training and testing, with further details in Subsection 4.2 and (Rasp et al., 2024). Moreover, we incorporate observation locations extracted from the real-time weather satellite track into the NWP experiment with higher spatial-resolution in Subsection 4.3.
@@
We compare our method against several baseline approaches: (1) 3D-Var with a known observation model, (2) a model-based 4D-Var algorithm that assumes known dynamical and observation models, (3) a learned inverse observation model with the known dynamical model, as proposed by (Frerix et al., 2021). In the two NWP problems, we include another three baselines: Latent 3D-Var and Latent 4D-Var (Cheng et al., 2024), which perform variational data assimilation in a latent space learned via an autoencoder, with Latent 4D-Var also modeling latent-space dynamics; and Fengwu 4D-Var (Xiao et al., 2024), an AI-embedded 4D-Var framework that integrates an advanced neural weather forecasting model into the DA process. These competitive baselines cover both operational Var-DA and ML-hybrid Var-DA methods.
Figure 2: Comparison of distribution of NRMSE (%) across different atmospheric variables (z500, t850, q700, u850, v850) for Latent 3D-Var, Latent 4D-Var, Fengwu 4D-Var, and Tensor-Var. The rightmost bar plot shows evaluation times and error bars indicate the standard deviation for evaluation time.
@@
In each experiment, we measure the quality of assimilation using the Normalized Root Mean Square Error (NRMSE) $\sqrt{\frac{1}{T}\sum_{t=0}^{T}\|\hat{s}_{t}-s_{t}\|^{2}}/(s_{\text{max}}-s_{\text{min}})$ over the assimilation window, where $s_{\text{max}},s_{\text{min}}$ are the maximum and minimum state values in the training dataset. These results and the average evaluation time for each algorithm are reported in Table 2. All metrics are evaluated 20 times with different initial conditions, reporting the mean and standard deviation. For Tensor-Var, the history length $m$ is selected using a cross-validation approach as an ablation study in Subsection 4.4, and the objective function in the feature space is minimized using quadratic programming, implemented via CVXPY (Diamond & Boyd, 2016). The baselines use the L-BFGS method (Nocedal & Wright, 2006) with 10 history vectors for the Hessian approximation. For 4D-Var baselines, we consider two cases (with and without adjoint models). The background state $s_{b}$ is set to the mean state of the training set.
As shown in Table 2, our method consistently outperforms the other baseline methods in all metrics. In all four tasks, Tensor-Var achieves the lowest mean and standard deviation of NRMSE for assimilation accuracy, demonstrating strong generalization from the Lorenz-96 systems to the more complex KS systems. The better performances are attributed to the linearization of the dynamics and the history-augmented inverse observation operator, which makes the 4D-Var optimization convex with more reliable convergence to the global optimum in the feature space. The larger NRMSE errors observed in the model-based 3D-Var and 4D-Var methods are due to the fact that incomplete observations lead to uncontrolled errors in the unobserved state dimensions. Although the ML-hybrid 4D-Var method (Frerix et al., 2021) employs a learned inverse observation model, it struggles with generalization due to the incomplete observations, leading to poor performance on test data. By incorporating historical information, our approach effectively controls estimation errors in the unobserved state dimensions. This improvement is demonstrated in Figures 7 and 8 in Appendix E.1.2, which qualitatively compares the assimilation performance in the Lorenz-96 and KS systems.
Figure 3: Comparison of assimilation NRMSE (%) over a 7-day horizon for five atmospheric variables from Latent 3D-Var, Latent 4D-Var, Fengwu 4D-Var, and Tensor-Var. Each time-step represents a 6-hour interval.
In terms of computational efficiency, the linear structure and convexity of the Tensor-Var enable a linear optimization convergence rate and much lower computational complexity, as shown in Table 1. Our results confirm these improvements by showing lower iteration times and overall evaluation times compared to other baselines. Moreover, the computational cost of feature mapping is negligible relative to the 4D-Var optimization, and the lowest NRMSE validates consistency of the optimization across both the original and feature spaces.